# Impact of Charge Transfer Inefficiency on transit light-curves

### v1

@@ Abstract
Context. PLATO is designed to detect Earth-sized exoplanets orbiting solar-type stars and to measure their radii (relative to the star radii) with an accuracy better than $2\%$ via the transit method. Charge transfer inefficiency (CTI), a by-product of radiation damage to charge-coupled devices (CCD), can jeopardise this accuracy constraint and therefore must be corrected to reach scientific requirements.
Aims. We assessed and quantified the impact of CTI on transit depth measurements. Our objective was to motivate the need for CTI correction and to develop a correction strategy that restores CTI-biased transit depths with an acceptable residual within the accuracy budget.
Methods. Using a calibration dataset generated with PLATOSim to simulate a realistic stellar field, we modelled the parallel overscan signal as a sum of exponential decays and used least-squares fitting to infer the number of trap species and initial estimates for the release times $\tau_{\mathrm{r,k}}$. Smearing was then modelled with an exponential-plus-constant function and removed on a column-wise basis. We modelled the spatial variation in the trap density with a quadratic polynomial function of the radial distance from the centre of the focal plane. The polynomial coefficients $a_{p,k}$ of this model, the well-fill power index $\beta$, and the release times $\tau_{\mathrm{r,k}}$ were subsequently adjusted via an iterative application of the Extended pixel Edge Response (EPER) method combined with a CTI correction algorithm. This yielded the final calibration model that underpins our correction strategy.
Results. In the worst-case scenario (8-year mission, high CTI impact zone), we found that CTI induced a bias of approximately $4\%$ in the measured transit depth. The polynomial coefficients from our trap density model were then used to correct the CTI-affected transit depths. Our correction reduced the bias to a residual of $0.06\%$, comfortably within PLATO’s accuracy requirements.
@@ 1 Introduction
PLATO (PLAnetary Transits and Oscillations of stars) is ESA’s M3 mission scheduled for launch in late 2026. PLATO aims to detect and describe exoplanets and conduct asteroseismological studies of the host and observed stars. Specifically, it will identify Earth-sized exoplanets (radius ¡ $2R_{Earth}$), orbiting F5-K7 dwarf and sub-giant stars (magnitude $\lesssim\penalty 10000\ 11$), with a particular focus on terrestrial planets located in the habitable zone of solar-type stars. Thus, it will offer critical insights about planet formation, existence of comparable solar systems and potentially habitable planets (Rauer et al., 2025).
@@
PLATO will conduct high precision photometric observations to identify these planets via the transit method. According to PLATO’s accuracy requirements, the planet-to-star radius ratio for an Earth-sized planet orbiting a G0V star of $m_{V}=10$ (goal $m_{V}=11$) must be determined with an accuracy better than $2\%$ (PLATO Science Working Team, 2019).
@@
This noise budget has to be shared among different sources of systematic error. However, as we will show in the present work, meeting these precision requirements will be challenging because the space environment can damage the charge-coupled devices (CCD) that enable the observations. The payload’s exposure to radiation will induce defects in the CCD, creating charge-trapping sites that intermittently capture and release electrons. This deferred charge release causes flux loss outside the aperture mask and degrades photometric observations. This phenomenon, known as charge transfer inefficiency (CTI), affects the quality of light-curves used to extract transit depths and, consequently, introduces measurement errors. While PLATO’s noise budget does not explicitly allocate CTI, the consortium’s internal reference value for sky background noise ($0.4\%$) serves as the benchmark threshold for CTI mitigation.
@@
Modelling and assessing the impact of CTI on astrophysical applications has been carried out for missions such as Euclid and FLEX (scheduled for launch in 2026) (Israel et al., 2015; Bernard et al., 2021). For Euclid’s CCD detectors, a CTI correction pipeline called ArCTIC, built on the Pyxel framework (Arko et al., 2022), has been developed (Kelman et al., 2022, 2024). While these existing methods rely on robust frameworks, they do not account for the spatial variation in trap density across the detector for each species. Ignoring such spatial variations may bias correction accuracy in real missions. Our work incorporates the spatially varying nature of trap density, which is anticipated for PLATO. To our knowledge, this is the first quantitative demonstration of CTI impacting exoplanet transit measurements. Although our study focuses on PLATO, the calibration and correction strategy presented here is broadly applicable to any detector system affected by CTI, providing a general framework for restoring CTI-impacted transit depths and preserving photometric accuracy.
@@
To motivate the need for CTI mitigation, we quantify the CTI-induced bias in the transit depth measurement of an Earth-sized exoplanet transiting a solar-type star in the context of PLATO’s accuracy requirements (Sect. 2). We present our CTI calibration strategy in Sect. 3. CTI is modelled using the Charge Density Model (CDM; Short et al. 2013; Massey et al. 2014), based on the Shockley–Read–Hall theory (Shockley & Read, 1952; Hall, 1952; Hardy et al., 1998). The CDM depends on the mean trap density ($n_{\mathrm{t}}$), release time ($\tau_{\mathrm{r}}$), well-fill power index ($\beta$), capture cross-section ($\sigma$), and operating temperature. To extend CDM beyond its uniform mean trap density assumption, we adopt a trap density distribution, incorporating an axisymmetric radial dependence inferred from the radiation map in OHB System AG (2024). The calibration proceeds in three stages: (1) initial iterative estimates of trap release times, (2) a non-iterative smearing correction via least squares, and (3) an iterative fit of the polynomial coefficients of the trap density model, well-fill power-law index $\beta$, and the release times via an algorithm involving the Extended pixel Edge Response (EPER) method combined with the Massey et al. (2014) CTI correction algorithm.
@@
Section 4 outlines our CTI correction strategy, adapted from the iterative inversion algorithm of Massey et al. (2014) and relying on the CTI parameters obtained from our calibration procedure. We evaluate the performance of this correction scheme, demonstrating its ability to recover CTI-affected transit depths toward their true values. Finally, Sect. 5 summarises the findings, highlights the limitations of the calibration–correction framework, and discusses the assumptions underlying our approach.
@@ 2.1 The PLATO instrument and the photometric measurements
The PLATO payload consists of 26 cameras: 24 “normal” cameras operating in white light for high-precision photometry, and 2 “fast” cameras providing colour information in the blue (505–700 nm) and red (665–1000 nm) bands (Marchiori et al., 2019). Each camera hosts four CCD detectors at its focal plane. About $80\%$ of PLATO targets will have on-board photometry extracted using fixed aperture masks. Given the limited on-board CPU resources, CTI correction at the pixel level is not feasible in this mode. For the remaining targets, which largely constitute the primary sample (S1), small $6\times 6$ pixel “imagettes” will be transmitted to ground at a cadence of 2525\,s for the normal cameras and 2.52.5\,s for the fast cameras. The photometry of these targets will be extracted on ground. In this work, we therefore designed and validated a CTI correction strategy specifically for the imagettes.
@@ 2.2 Basic characteristics of CTI
CTI is a by-product of prolonged exposure of CCD to radiation in space, from sources such as cosmic rays and solar particles. Radiation-induced defects create trapping sites in the silicon lattice that capture and later release charges during transfer stage in the CCD. During its in-flight operation, the instruments’ prolonged exposure to radiation will increase the number of charge trapping sites and thus, intensify CTI. This will impact the measured signal, causing charge trailing along the CCD columns (Fig. 1). This delayed release of electrons causes redistribution of flux from the source, leading to distorted image and potential photometric biases.
Figure 1: Effect of CTI on the recorded signal in the parallel transfer direction (readout proceeds from left to right). The solid curve shows a CTI-free signal profile, while the dashed curve illustrates the same profile after CTI degradation. Deferred charge release reduces the peak amplitude and produces a trailing toward increasing TDI lines (earlier pixels in the transfer path). Credit: Short et al. (2013).
CTI is a serious concern for PLATO, which relies on photometric time series to detect transits and identify exoplanets, as the charge-release timescale is of the order of $10^{-4}\,\mathrm{s}$ for the fastest traps and $10^{-1}\,\mathrm{s}$ for the slowest traps. These values are – for the majority of the trap species – significantly longer (particularly for the slowest traps) than the transfer cadence of about $900\penalty 10000\ \mu\mathrm{s}$, posing a risk to the robustness of the observations.
@@ 2.3 CTI modelling
Various strategies exist to simulate parallel CTI effects, differing mainly in their treatment of trap occupancy and integration. In this work, we used the Charge Density Model (CDM), which models charge capture and release probabilistically and governs the interaction between traps and the electron cloud (see e.g. and reference therein Short et al., 2013; Massey et al., 2014). Specifically, following the formalism of Short et al. (2013), the probability of trap capture depends on the local electron density and is controlled by the parameter $\beta$, which defines how the electron cloud volume scales with the number of electrons. A larger electron cloud increases trap accessibility, and this dependence is expressed as
@@
$V_{\mathrm{c}}$ volume of the charge cloud in a pixel;
$V_{\mathrm{g}}$ maximum geometrical volume a pixel can hold;
$N_{\mathrm{e}}$ number of electrons in the pixel;
$\mathrm{FWC}$ full-well capacity of the pixel (maximum electron load);
$\beta$ well-fill power-law index ($0\leq\beta\leq 1$ in practice).
CTI can be modelled using the run_cdm_parallel routine in the Pyxel [^1] https://esa.gitlab.io/pyxel/ framework (Arko et al., 2022). Based on Short et al. (2013) formulation, this routine requires as inputs: the well-fill power index $\beta$, the trap release times $\tau_{\mathrm{r}}$, the trap densities $n_{\mathrm{t}}$, and capture cross-sections $\sigma$, which control the interaction probability between charges and traps for each involved trap species. Short et al. (2013) formalism assumes uniform trap densities. In practice, this is not the case as we describe in Sect. 2.4. However, as explained below, it is possible to generalize this formalism in order to account for an arbitrary distribution of traps.
For a uniform trap distribution, the number of traps (for a given trap species $k$) encountered by a charge packet during transfer from row index $i$ to the register is given by $n_{t,k}\,y$, where $n_{t,k}$ is the (constant) trap density and $y=i+1$ denotes the number of transfers (i.e., the distance from row index $i$ to the register).
When the trap distribution varies spatially, the term $n_{t,k}\,y$ is replaced by $cnt_{k}[i,j]$, which represents the cumulative number of traps crossed by a charge packet during its transfer from row $i$ down to the register, for a given column $j$. We extended the CDM accordingly by introducing run_cdm_parallel_cumul (see Python implementation in Appendix A). Unlike the original routine, which assumes a constant trap density per pixel, our adaptation accounts for the cumulative trap content encountered during vertical transfer. This captures the integrated effect of spatial variations in the trap distribution.
Trap species parameters from Prod’homme et al. (2016), including the well-fill power index $\beta$, release times $\tau_{r,k}$, capture cross-sections $\sigma_{k}$, and reference mean trap densities $n_{t,k}^{\mathrm{ref}}$, are listed in Table 1. In this study, we adopted these reference values: $\beta=0.37$, species-specific $\sigma_{k}$ and $\tau_{r,k}$. The trap densities $n_{t,k}$ are subsequently employed to construct the cumulative trap distribution, as described in Sect. 2.4.
Table 1: Reference trap species parameters from Prod’homme et al. (2016) for a fluence of $13\times 10^{9}\,\mathrm{cm}^{-2}$. The well-fill power index is fixed at $\beta=0.37$.
@@ 2.4 Radiation analysis
The PLATO radiation analysis by OHB System AG (2024) reports the total non-ionising dose (TNID) for Camera 1, which comprises four CCD, after 6.5 years of exposure. The radiation map, displayed in Fig. 2, indicates that TNID—and therefore trap density—increases approximately with the square of the radial distance from the centre, implying a first-order axisymmetric distribution across the CCD.
Figure 2: Radiation map of Camera 1, composed of four CCD, showing the radial dependence of trap density. The colour bar indicates the TNID level (trap density) to which each pixel is exposed after 6.5 years. Adapted from OHB System AG (2024).
To reproduce the radiation map shown in Fig. 2, we generated a grid of coordinates across the CCD and computed the normalised radial distance $R_{\mathrm{norm}}$ using
@@
The TNID distribution is approximately modelled as a quadratic function of the normalised radial distance $R_{\mathrm{norm}}$
@@
varying between a minimum dose $D_{\mathrm{min}}$ at the radial origin and a maximum dose $D_{\mathrm{max}}$ at the edge. Fig. 3 compares the radiation map derived from Eq. 3 with the reference distribution in Fig. 2, showing close agreement for CCD #1 under identical TNID conditions. The residual departure from our axisymmetric model is discussed in Sect. 5.
@@
For a 6.5-year mission, the maximum TNID reported by OHB System AG (2024) corresponds to a total fluence of $5\times 10^{9}\,\mathrm{cm}^{-2}$. Assuming a constant radiation rate, the annual fluence is computed by linear scaling
@@
The TNID for 4-year (nominal) and 8-year (extended) duration is then obtained by multiplying $\Phi_{\mathrm{year}}$ by the respective mission duration. The reference mean trap densities $n_{t,k}^{\mathrm{ref}}$ in Table 1 correspond to a TNID of $13\times 10^{9}\,\mathrm{cm}^{-2}$ (Prod’homme et al., 2016). To account for mission-specific doses, these values are scaled linearly as
$n_{t,k}^{\mathrm{scaled}}=\left(\frac{\Phi_{\mathrm{year}}\times\mathrm{T}}{13\times 10^{9}}\right)n_{t,k}^{\mathrm{ref}}\,$ (5)
where T is the desired mission duration in years. The scaled trap densities are then distributed across the CCD according to an axisymmetric radial model, that is
@@
A cumulative sum along the row axis yields a three-dimensional array representing the cnt parameter, that is
@@
This latter quantity serves as input for the CTI simulations in subsequent steps.

### v2

# Impact of charge transfer inefficiency on transit light curves

@@ Abstract
Context. PLATO is designed to detect Earth-sized exoplanets orbiting solar-type stars and to measure their radii (relative to the star radii) with an accuracy better than $2\%$ via the transit method. Charge transfer inefficiency (CTI), a by-product of radiation damage to charge-coupled devices (CCDs), can jeopardise this accuracy constraint and therefore must be corrected to reach scientific requirements.
Aims. We assessed and quantified the impact of CTI on transit depth measurements. Our objective was to demonstrate the need for CTI correction and to develop a correction strategy that restores CTI-biased transit depths with an acceptable residual within the accuracy budget.
Methods. Using a calibration dataset generated with PLATOSim to simulate a realistic stellar field, we modelled the parallel overscan signal as the sum of exponential decays and used least-squares fitting to infer the number of trap species and initial estimates for the release times ($\tau_{\mathrm{r,k}}$). Smearing was then modelled with an exponential-plus-constant function and removed on a column-wise basis. We modelled the spatial variation in the trap density with a quadratic polynomial function of the radial distance from the centre of the focal plane. The polynomial coefficients ($a_{p,k}$) of this model, the well-fill power index ($\beta$), and the release times ($\tau_{\mathrm{r,k}}$) were subsequently adjusted via an iterative application of the extended pixel edge response method combined with a CTI correction algorithm. This yielded the final calibration model that underpins our correction strategy.
Results. In the worst-case scenario (8-year mission, high CTI impact zone), we found that CTI induced a bias of approximately $4\%$ in the measured transit depth. The polynomial coefficients from our trap density model were then used to correct the CTI-affected transit depths. Our correction reduced the bias to a residual of $0.06\%$, which is comfortably within PLATO’s accuracy requirements.
@@ 1 Introduction
The PLAnetary Transits and Oscillations of stars (PLATO) is the European Space Agency’s M3 mission and is scheduled for launch in late 2026. The goal of the PLATO mission is to detect and describe exoplanets and conduct asteroseismological studies of host and observed stars. Specifically, it will identify Earth-sized exoplanets (radius ¡ $2R_{Earth}$) orbiting F5-K7 dwarf and sub-giant stars (magnitude $\lesssim\penalty 10000\ 11$), with a particular focus on terrestrial planets located in the habitable zone of solar-type stars. Thus, it will offer critical insights into planet formation and the existence of comparable solar systems and potentially habitable planets (Rauer et al., 2025).
@@
PLATO will conduct high-precision photometric observations to identify these planets via the transit method. According to PLATO’s accuracy requirements, the planet-to-star radius ratio for an Earth-sized planet orbiting a G0V star of $m_{V}=10$ (goal $m_{V}=11$) must be determined with an accuracy better than $2\%$ (PLATO Science Working Team, 2019).
@@
This noise budget has to be shared among different sources of systematic error. However, as we show in the present work, meeting these precision requirements will be challenging because the space environment can damage the charge-coupled devices (CCDs) that enable the observations. The payload’s exposure to radiation will induce defects in the CCDs, creating charge-trapping sites that intermittently capture and release electrons. This deferred charge release causes flux loss outside the aperture mask and degrades photometric observations. This phenomenon, known as charge transfer inefficiency (CTI), affects the quality of light curves used to extract transit depths and, consequently, introduces measurement errors. While PLATO’s noise budget does not explicitly allocate CTI, the consortium’s internal reference value for sky background noise ($0.4\%$) serves as the benchmark threshold for CTI mitigation.
@@
Modelling and assessing the impact of CTI on astrophysical applications has been carried out for missions such as Euclid and the FLuorescence EXplorer (FLEX; scheduled for launch in 2026; Israel et al. 2015; Bernard et al. 2021). For Euclid’s CCD detectors, a CTI correction pipeline called ArCTIC, built on the Pyxel framework (Arko et al., 2022), has been developed (Kelman et al., 2022, 2024). While these existing methods rely on robust frameworks, they do not account for the spatial variation in trap density across the detector for each species. Ignoring such spatial variations may bias correction accuracy in real missions. Our work incorporates the spatially varying nature of trap density, which is anticipated for PLATO. To our knowledge, this is the first quantitative demonstration of CTI impacting exoplanet transit measurements. Although our study focuses on PLATO, the calibration and correction strategy presented here is broadly applicable to any detector system affected by CTI, providing a general framework for restoring CTI-impacted transit depths and preserving photometric accuracy.
@@
To demonstrate the need for CTI mitigation, we quantified the CTI-induced bias in the transit depth measurement of an Earth-sized exoplanet transiting a solar-type star in the context of PLATO’s accuracy requirements (Sect. 2). We present our CTI calibration strategy in Sect. 3. CTI is modelled using the charge density model (CDM; Short et al. 2013; Massey et al. 2014), which is based on the Shockley–Read–Hall theory (Shockley and Read, 1952; Hall, 1952; Hardy et al., 1998). The CDM depends on the mean trap density ($n_{\mathrm{t}}$), release time ($\tau_{\mathrm{r}}$), well-fill power index ($\beta$), capture cross-section ($\sigma$), and operating temperature. To extend CDM beyond its uniform mean trap density assumption, we adopted a trap density distribution, incorporating an axisymmetric radial dependence inferred from the radiation map in OHB System AG (2024). The calibration proceeds in three stages: (1) initial iterative estimates of trap release times, (2) a non-iterative smearing correction via least squares, and (3) an iterative fit of the polynomial coefficients of the trap density model, well-fill power-law index ($\beta$), and the release times via an algorithm involving the extended pixel edge response (EPER) method combined with the Massey et al. (2014) CTI correction algorithm.
@@
Section 4 outlines our CTI correction strategy, which is adapted from the iterative inversion algorithm of Massey et al. (2014) and relies on the CTI parameters obtained from our calibration procedure. We evaluate the performance of this correction scheme, demonstrating its ability to recover CTI-affected transit depths toward their true values. Finally, Sect. 5 summarises our findings, highlights the limitations of the calibration-correction framework, and discusses the assumptions underlying our approach.
@@ 2.1 The PLATO instrument and the photometric measurements
The PLATO payload consists of 26 cameras: 24 ‘normal’ cameras operating in white light for high-precision photometry, and 2 ‘fast’ cameras providing colour information in the blue (505–700 nm) and red (665–1000 nm) bands (Marchiori et al., 2019). Each camera hosts four CCD detectors at its focal plane. About $80\%$ of PLATO targets will have on-board photometry extracted using fixed aperture masks. Given the limited on-board CPU resources, CTI correction at the pixel level is not feasible in this mode. For the remaining targets, which largely constitute the primary sample (S1), small $6\times 6$ pixel ‘imagettes’ will be transmitted to ground at a cadence of 2525\,s for the normal cameras and 2.52.5\,s for the fast cameras. The photometry of these targets will be extracted on ground. In this work, we therefore designed and validated a CTI correction strategy specifically for the imagettes.
@@ 2.2 Basic characteristics of CTI
Charge transfer inefficiency is a by-product of prolonged exposure of CCDs to radiation in space, from sources such as cosmic rays and solar particles. Radiation-induced defects create trapping sites in the silicon lattice that capture and later release charges during transfer stage in the CCDs. During its in-flight operation, the instruments’ prolonged exposure to radiation will increase the number of charge trapping sites and thus, intensify CTI. This will impact the measured signal, causing charge trailing along the CCD columns (Fig. 1). This delayed release of electrons causes redistribution of flux from the source, leading to distorted image and potential photometric biases.
Figure 1: Effect of CTI on the recorded signal in the parallel transfer direction (readout proceeds from left to right). The solid curve shows a CTI-free signal profile, while the dashed curve illustrates the same profile after CTI degradation. Deferred charge release reduces the peak amplitude and produces a trailing towards increasing time delay integration (TDI) lines (earlier pixels in the transfer path). Credit: Short et al. (2013).
Charge transfer inefficiency is a serious concern for PLATO, which relies on photometric time series to detect transits and identify exoplanets, as the charge-release timescale is of the order of $10^{-4}\,\mathrm{s}$ for the fastest traps and $10^{-1}\,\mathrm{s}$ for the slowest traps. These values are – for the majority of the trap species – significantly longer (particularly for the slowest traps) than the transfer cadence of about $900\penalty 10000\ \mu\mathrm{s}$, posing a risk to the robustness of the observations.
@@ 2.3 CTI modelling
Various strategies exist to simulate parallel CTI effects, differing mainly in their treatment of trap occupancy and integration. In this work, we used the CDM, which models charge capture and release probabilistically and governs the interaction between traps and the electron cloud (see e.g. Short et al., 2013; Massey et al., 2014, and reference therein). Specifically, following the formalism of Short et al. (2013), the probability of trap capture depends on the local electron density and is controlled by the parameter $\beta$, which defines how the electron cloud volume scales with the number of electrons. A larger electron cloud increases trap accessibility, and this dependence is expressed as
@@
where $V_{\mathrm{c}}$ is volume of the charge cloud in a pixel; $V_{\mathrm{g}}$ is maximum geometrical volume a pixel can hold; $N_{\mathrm{e}}$ is the number of electrons in the pixel; $\mathrm{FWC}$ is full-well capacity of the pixel (the maximum electron load); and $\beta$ is the well-fill power-law index ($0\leq\beta\leq 1$ in practice).
Charge transfer inefficiency can be modelled using the run_cdm_parallel routine in the Pyxel [^1] https://esa.gitlab.io/pyxel/ framework (Arko et al., 2022). Based on the Short et al. (2013) formulation, this routine requires as inputs: the well-fill power index ($\beta$), the trap release times ($\tau_{\mathrm{r}}$), the trap densities ($n_{\mathrm{t}}$), and capture cross-sections ($\sigma$), which control the probability of interactions between charges and traps for each involved trap species. Short et al. (2013) formalism assumes uniform trap densities. In practice, this is not the case as we describe in Sect. 2.4. However, as explained below, it is possible to generalise this formalism in order to account for an arbitrary distribution of traps.
For a uniform trap distribution, the number of traps (for a given trap species $k$) encountered by a charge packet during transfer from row index $i$ to the register is given by $n_{t,k}\,y$, where $n_{t,k}$ is the (constant) trap density and $y=i+1$ denotes the number of transfers (i.e. the distance from row index $i$ to the register).
When the trap distribution varies spatially, the term $n_{t,k}\,y$ is replaced by $cnt_{k}[i,j]$, which represents the cumulative number of traps crossed by a charge packet during its transfer from row $i$ down to the register, for a given column $j$. We extended the CDM accordingly by introducing run_cdm_parallel_cumul (see the Python implementation in plato-cti-correction). Unlike the original routine, which assumes a constant trap density per pixel, our adaptation accounts for the cumulative trap content encountered during vertical transfer. This captures the integrated effect of spatial variations in the trap distribution.
Trap species parameters from Prod’homme et al. (2016), including the well-fill power index ($\beta$), release times ($\tau_{r,k}$), capture cross-sections ($\sigma_{k}$), and reference mean trap densities ($n_{t,k}^{\mathrm{ref}}$), are listed in Table 1. In this study, we adopted these reference values: $\beta=0.37$, species-specific $\sigma_{k}$ and $\tau_{r,k}$. The trap densities ($n_{t,k}$) are subsequently employed to construct the cumulative trap distribution, as described in Sect. 2.4.
Table 1: Reference trap species parameters. [^2] Parameters are from Prod’homme et al. (2016) for a fluence of $13\times 10^{9}\,\mathrm{cm}^{-2}$. The well-fill power index is fixed at $\beta=0.37$.
@@ 2.4 Radiation analysis
The PLATO radiation analysis by OHB System AG (2024) reports the total non-ionising dose (TNID) for Camera 1, which comprises four CCDs, after 6.5 years of exposure. The radiation map, displayed in Fig. 2, indicates that TNID—and therefore trap density—increases approximately with the square of the radial distance from the centre, implying a first-order axisymmetric distribution across the CCD.
Figure 2: Radiation map of Camera 1, composed of four CCDs, showing the radial dependence of the trap density. The colour bar indicates the TNID level (trap density) to which each pixel is exposed after 6.5 years. Adapted from OHB System AG (2024).
To reproduce the radiation map shown in Fig. 2, we generated a grid of coordinates across the CCD and computed the normalised radial distance ($R_{\mathrm{norm}}$) using
@@
The TNID distribution is approximately modelled as a quadratic function of the normalised radial distance ($R_{\mathrm{norm}}$):
@@
varying between a minimum dose ($D_{\mathrm{min}}$) at the radial origin and a maximum dose ($D_{\mathrm{max}}$) at the edge. Figure 3 compares the radiation map derived from Eq. 3 with the reference distribution in Fig. 2, showing close agreement for CCD #1 under identical TNID conditions. The residual departure from our axisymmetric model is discussed in Sect. 5.
@@
For a 6.5-year mission, the maximum TNID reported by OHB System AG (2024) corresponds to a total fluence of $5\times 10^{9}\,\mathrm{cm}^{-2}$. Assuming a constant radiation rate, the annual fluence is computed via linear scaling:
@@
The TNID for 4-year (nominal) and 8-year (extended) duration is then obtained by multiplying $\Phi_{\mathrm{year}}$ by the respective mission duration. The reference mean trap densities ($n_{t,k}^{\mathrm{ref}}$) in Table 1 correspond to a TNID of $13\times 10^{9}\,\mathrm{cm}^{-2}$ (Prod’homme et al., 2016). To account for mission-specific doses, these values are scaled linearly as
$n_{t,k}^{\mathrm{scaled}}=\left(\frac{\Phi_{\mathrm{year}}\times\mathrm{T}}{13\times 10^{9}}\right)n_{t,k}^{\mathrm{ref}}\,$, (5)
where T is the desired mission duration in years. The scaled trap densities are then distributed across the CCD according to an axisymmetric radial model, that is,
@@
A cumulative sum along the row axis yields a three-dimensional array representing the cnt parameter:
@@
This quantity serves as input for the CTI simulations in subsequent steps.